OpenStudy (anonymous):

You have just graduated and are offered a job. The starting salary is $25,000 per year and you are told that you will receive either: (A) a $1000/yr raise each year [a Geometric Sequence] or (B) a 3% raise each year [an Arithmetic Sequence] . You are to determine which of the two choices would give you a better salary after 5 years.

OpenStudy (dumbcow):

its actually the other way around adding 1000 each year is Arithmetic adding 3% is like multiplying by 1.03 each year (Geometric) For the Arithmetic: a1 = 25,000, d=1000 a5 = 25,000 + 1000(5-1) = 29,000 For geometric: a1 =25,000, r = 1.03 a5 = 25,000(1.03)^(5-1) = 28,137 So in short term, getting 1000 each year is better

OpenStudy (anonymous):

lol good thing i deleted my answer to avoid embarassment

OpenStudy (anonymous):

DUMB cow is right

OpenStudy (dumbcow):

aww i didn't even get a chance to look at it :)

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Dumb cow ur amazing thanks soooo much:)